A force of 250 N exerted at an angle of 30° above the horizontal is needed to keep a 90 kg block moving along a

cement floor. Find the coefficient of kinetic friction.


1
Expert's answer
2022-04-07T13:09:02-0400
"\mu=\frac{F\cos\theta}{mg-F\sin\theta}"

"\mu=\frac{250*\cos30^\circ}{90*9.8-250*\sin30^\circ}=0.286"